OpenStudy (anonymous):

simplify the sum of (w^2+2w-24)/(w^2+w-30)+8/(w-5)

OpenStudy (anonymous):

\[\frac{w^2+2w-24}{(w-5)(w+6)}+\frac{8}{w-5}\] \[=\frac{w^2+2w-24+8(w+6)}{(w-5)(w+6)}\] then algebra in the numerator distribute and combine like terms